OpenStudy (anonymous):

HELPPP. Does anyone know what formula i would use for this ? 5. Jackie deposited $5 into a checking account in February. For each month following, the deposit amount was doubled. How much money was deposited in the checking account in the month of August?

OpenStudy (amistre64):

f5 m10 a20 m40 j80 j160 a ...

OpenStudy (anonymous):

huh ?

OpenStudy (amistre64):

Bn = B0*2^(n-1) but then youd need to adjust for feb = 0 and calculate the n for august

OpenStudy (amistre64):

you double the amount each month ... its not that hard of a brute math that a formula is needed.

OpenStudy (anonymous):

i just need to know the name of the formula .

OpenStudy (amistre64):

a name eh ... cant say that i know their names in particular

OpenStudy (amistre64):

geometric comes to mind .... but not sure if thats what is needed

OpenStudy (anonymous):

i thought geometric sequence but im not sure.

OpenStudy (amistre64):

feb = 5 mar = 5*2 apr = 5*2*2 may = 5*2*2*2 its geometric

OpenStudy (anonymous):

thanks . i already had the answer and worked it out. i just didnt remember what formula i used.

OpenStudy (amistre64):

:)